Can someone help me on this problem?
I need to add:
2 + (-11) + 14 + (-7)
Thanks,
Sher

Found 2 solutions by AnlytcPhil, pwt:
Answer by AnlytcPhil(1806)

2 + (-11) + 14 + (-7) 

Add the first two, 2 + (-11).  The signs are different, so
you subtract 11-2 and get 9, then since 11 is larger and it has
a - sign, you put a negative sign onto 9, and get -9. So you
replace the 2 + (-11) by -9. That is,

2 + (-11) + 14 + (-7)  becomes

       -9 + 14 + (-7)

Next add the first two, -9 + 14.  The signs are different, so
you subtract 14-9 and get 5, then since 14 is larger and it has
a + sign, you put a positive sign onto 5, and get +5. But then
you can just leave the + sign understood.  So you replace the
-9 + 14 by 5. That is,

       -9 + 14 + (-7)   becomes

             5 + (-7)

Finally add the 5 + (-7).  The signs are different, so
you subtract 7-5 and get 2, then since 7 is larger and it has
a - sign, you put a negative sign onto 2, and get -2. So the
answer is

                -2

Edwin


Answer by pwt(2)

2+(-11)+14+(-7)=2+14-11-7=16-18=(-2)